The displacement x of a particle is given by x = A sin (ωt + θ), the time at which the displacement is maximum is. can you please explain how to solve it??

Options:

 

Grade:12th pass

1 Answers

Krishan Pareek
6 years ago
Here the quantity A is constant and therefore x is maximum when sin(wt+€) is maximum. sin is max when angle is π/2wt+€= π/2wt= π/2 -€t= (π/2-€)/w